A project has the following cash flows:

Year Cash Flow
0 –$ 17,200
1 7,900
2 9,200
3 7,700

   

a.

What is the NPV at a discount rate of zero percent? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

b. What is the NPV at a discount rate of 12 percent? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
c. What is the NPV at a discount rate of 22 percent? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
d. What is the NPV at a discount rate of 28 percent? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Please show how you get Present value at 12%, 22%, & 28% Please!
Answer #1

a.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=7900+9200+7700=24800

NPV=Present value of inflows-Present value of outflows

=24800-17200

=$7600

b.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=7900/1.12+9200/1.12^2+7700/1.12^3

=19868.46

NPV=Present value of inflows-Present value of outflows

=19868.46-17200

=$2668.46(Approx)

c.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=7900/1.22+9200/1.22^2+7700/1.22^3

=16896.99

NPV=Present value of inflows-Present value of outflows

=16896.99-17200

=$-303.01(Approx)(Negative)

d.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=7900/1.28+9200/1.28^2+7700/1.28^3

=15458.76

NPV=Present value of inflows-Present value of outflows

=15458.76-17200

=$-1741.24(Approx)(Negative)